Garnett Co. expects to purchase $180,000 of materials in July and $210,000 of materials in August. Three-fourths of all purchases are paid for in the month of purchase, and the other one-fourth are paid for in the month following the month of purchase. How much will August's cash disbursements for materials purchases be?
Answer #1

$202,500
Answer #2

180,000 x .25 + 210,000 x .75 = 45,000 + 157,500 = 202,500
Answer #3

July August September

Purchases 135,000 157,500   

45,000 52,500

Total disbursement for august $202,500